Platinum Water Health is evaluating the laundromat project. During year 1, the laundromat project is expected to have relevant revenue of 666,900 dollars, relevant variable costs of 221,100 dollars, and relevant depreciation of 65,800 dollars. In addition, Platinum Water Health would have one source of fixed costs associated with the laundromat project. Platinum Water Health just signed a deal with Orange Valley Media to develop an advertising campaign for use in the project. The terms of the deal require Platinum Water Health to pay Orange Valley Media either 109,100 dollars in 1 year if the project is pursued or 149,200 dollars in 1 year if the project is not pursued. Relevant net income for the laundromat project in year 1 is expected to be 220,510 dollars. What is the tax rate expected to be in year 1?  Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Answer:

Assuming that the project will be pursued the income before tax is calculated as follows:

$666,900 Revenue $221,100 Costs $65,800 Depreciation Advertisning expense Total operating expenses Operating Income $109,100

We observe Income before tax = $270,900

Given that net income of year 1 = $220,510

Hence:

Tax = 270900 - 220510 = $50,390

Expected Tax rate in year 1 = 50390 / 270900 = .1860 or 18.60%

Expected Tax rate in year 1 = .1860